# Align-GRAG: Reasoning-Guided Dual Alignment for Graph Retrieval-Augmented Generation

### v2

@@ 5.6. Generalization Analysis
GraphTransformer GNP 0.8704 0.7391 0.8880 0.7696 0.8892 0.7557
G-Retriever 0.8705 0.7366 0.9115 0.7739 0.9014 0.7482
Align-GRAG 0.8992 0.7626 0.9241 0.7789 0.9321 0.7712
GAT GNP 0.9061 0.7291 0.8989 0.7676 0.9062 0.7521
G-Retriever 0.7960 0.7414 0.8953 0.7737 0.9111 0.7719
Align-GRAG 0.9151 0.7309 0.9151 0.7573 0.9214 0.7742
GCN GNP 0.7545 0.7298 0.8682 0.7564 0.8572 0.7495
G-Retriever 0.8592 0.7352 0.9007 0.7521 0.8512 0.7527
Align-GRAG 0.8574 0.7377 0.9152 0.7592 0.8596 0.7573
In this section, we evaluate whether Align-GRAG generalizes across both GNN architectures and LLM scales on ExplaGraphs (Accuracy) and WebQSP (Hit@1) datasets. Table 3 summarizes results for three GNN backbones—GT (GraphTransformer) (Shi et al., 2020), GAT (Veličković et al., 2018), and GCN (Kipf and Welling, 2016)—and three LLM families/sizes (Llama-2-7b-hf, Llama-2-13b-hf, Llama-3.1-8B), comparing against GNP and G-Retriever. We can observe that Align-GRAG attains the best score in 15/18 configurations, indicating robust generalization to diverse encoder–reasoner pairings.
What is more, we can find the strongest gains with GraphTransformer (GT). Using GT as the graph encoder yields the largest margins over GNP and G-Retriever on both datasets and across LLM sizes. This suggests GT’s capacity to model long-range dependencies synergizes with Align-GRAG’s dual alignment of nodes and representations, amplifying retrieval and reasoning quality.
Besides, benefits persist across attention-based GNNs. With GAT, Align-GRAG remains consistently competitive and often best, reinforcing that alignment mechanisms complement attention-driven neighborhood aggregation. Larger LLMs help, especially with GCN. Performance with 13B models generally surpasses 7B counterparts; the gap is most pronounced under the GCN backbone, where the LLM must compensate for weaker expressivity in the graph encoder. This indicates that larger LLMs better exploit aligned graph signals for multi-hop reasoning.
On the other hand, improvements hold on both ExplaGraphs and WebQSP datasets, demonstrating that Align-GRAG transfers well from explanation-centric graph tasks to question answering over knowledge graphs. Overall, Align-GRAG is architecture-agnostic yet particularly effective when paired with more expressive GNNs and larger LLMs, underscoring the importance of both high-capacity graph encoders and language models for learning fine-grained node-representation alignment.
5.7. Evaluation of Representation Alignment
@@
(b) Align with Summarization
@@
Figure 4. Representation Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we calculate the Cosine Similarity scores between the graph embeddings and the language embeddings of the query, summarization, and textualized graph, respectively, on test set. The graph embeddings include both the unaligned embeddings and the aligned embeddings (optimized using our contrastive loss). From the results in Figure 4, we observe that as the alignment degree increases, Cosine Similarity scores improve. For instance, similarity with summarization embeddings improves significantly as alignment progresses from 40 to 80 steps. For query and textualized graph embeddings, scores rise sharply between 0-4 steps. This indicates that training effectively reduces the representation gap between graph and language embeddings. After 8 steps, aligned embeddings consistently achieve higher similarity scores than unaligned across all text representations (query, summarization, and textualized graph), demonstrating the effectiveness of our approach. However, as shown in section 5.9.2, the alignment degree does not necessarily improve indefinitely. When it reaches a certain level, excessive alignment may over-alter the original graph information, compromising its accuracy.
5.8. Evaluating the Efficiency of Node Pruning
Table 4. Evaluation of Efficiency with and without node pruning. $\uparrow$ indicates results better than G-Retriever, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to G-Retriever. OOM means out of memory.
@@ 5.8. Evaluating the Efficiency of Node Pruning
No Pruning Non-Retriever 100626.28 (387.02%) OOM -
BM25 (Robertson et al., 2009) 2569.57 (99.72%) 17:24 min 0.4394
Contriever (Izacard et al., [n. d.]) 2593.28 (100.01%) 17:37 min 0.4429
G-Retriever (Bienstock et al., 1993) 2576.66 (100.00%) 17:28 min 0.4502
Node Pruning with
Different Degree
@@
In this section, we evaluate the efficiency improvements brought by node pruning in Aligner module. Specifically, we compare different retrieval methods in terms of average token consumption, average inference time, and Hit@1 scores on the test set of WebQSP. The w/ Aligner method is built on the initial retrieval results from PCST as the same as G-Retriever (He et al., 2024), where different seed nodes are selected to perform node pruning. The experiments were conducted in an inference-only setting. To fully utilize GPU resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 and Contriever methods retrieve triples and ensure that the token consumption is roughly similar to that of G-Retriever. From Table 4, we observe that when the number of seed nodes ($n_{\text{seed}}$) is set to 6, our Aligner method achieves performance comparable to G-Retriever while only utilizing 27.11% tokens. This demonstrates the method’s ability to effectively prune irrelevant knowledge while preserving the essential information. However, when $n_{\text{seed}}$ is too small (e.g., $n_{\text{seed}}=4$), performance drops significantly, suggesting that some critical information may be pruned during the process. On the other hand, increasing the number of seed nodes to 15 (utilizing 60.04% of the tokens) leads to performance that significantly surpasses G-Retriever. This highlights the strength of our method in efficiently selecting useful knowledge and confirms that, in longer-context scenarios, removing irrelevant information enhances overall results. Additionally, we find that BM25 and Contriever both perform worse than G-Retriever when using a similar number of tokens. This suggests that the PCST retrieval method is better suited for graph databases, as it more effectively captures the connections between knowledge.
5.9. Impact of Hyperparameters
@@
5.9.1. Impact of Seed Nodes
@@
Figure 5. Hyperparameters Analysis of the Number of seed nodes.
5.9.2. Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 6, we evaluate the Hit@1, F1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1, F1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the representation alignment is most effective, and training beyond this point can be detrimental.
Figure 6. Hyperparameters Analysis of Align degree.
5.9.3. Impact of Top K retrieval
@@
Figure 7. Effect of Top K retrieval.
5.10. Evaluating the Quality of Reasoning Chain
In this section, we conduct a fine-grained analysis of the quality of the reasoning chain by LLM summarization. This analysis is divided into two aspects. First, we evaluate Relevance and Faithfulness to assess whether the generated chains are aligned with the questions and answers. Second, we compare the QA performance of the naive-quality reasoning chain to highlight our method’s impact. What is more, a case study is used to gain a deeper understanding of the forms of summarization in Section 5.11.
5.10.1. Evaluation of Relevance and Faithfulness
To assess the quality of generated reasoning chain in terms of their relevance to question-answer context and their faithfulness to provided contextual information, we adopted LLM-as-a-judge framework (Gu et al., 2024). Specifically, we utilized GPT-4o (OpenAI, 2024), genimi-2.5-flash (Comanici et al., 2025) and human to perform the evaluation. A total of 100 samples were randomly selected for this analysis. The evaluation was conducted using a carefully designed prompt, as illustrated in Figure 10. The results of the experiment are presented in Figure 8. The analysis reveals that the summarized reasoning chains exhibit a high degree of relevance and faithfulness. Specifically, the summaries achieved relevance scores of 90%, 92%, and 92%, indicating that the majority of the generated content aligns well with the given question-answer context. Similarly, faithfulness scores of 87%, 89%, and 90% demonstrate that the summaries largely adhere to the factual information provided by the context, with minimal hallucination. Furthermore, the alignment between LLM evaluations and human evaluations indicates the reliability and rationality of LLM-based judge.
Figure 8. Evaluating Relevance and Faithfulness of Reasoning Chain.
5.10.2. Quantitative comparison
In this experiment, we evaluate the impact of reasoning quality on QA performance by comparing two approaches: our $\text{Reasoning}_{\text{Align-GRAG}}$, where both the question and answer are included during reasoning chain generation, and $\text{Reasoning}_{\text{Naive}}$, where only the question is used for reasoning chain generation. The results in Table 5 demonstrate that incorporating the answer during summary generation ($\text{Reasoning}_{\text{Align-GRAG}}$) consistently outperforms the naive approach ($\text{Reasoning}_{\text{Naive}}$) across all evaluation metrics of datasets. Specifically, for ExplaGraphs, the accuracy rises significantly from 0.8592 using a naive reasoning approach to 0.8992 when employing $\text{Reasoning}_{\text{Align-GRAG}}$. For SceneGraphs, the accuracy improves from 0.8404 to 0.8804, illustrating the advantage of our method in enhancing graph-based reasoning accuracy. For WebQSP, The F1 score of $\text{Reasoning}_{\text{Align-GRAG}}$ is 0.5445, which is higher than the 0.5209 achieved by $\text{Reasoning}_{\text{Naive}}$, indicating that our strategy improves the balance of precision and recall in QA tasks. Similarly, the Hit@1 value increases from 0.7444 to 0.7626, showing that our approach helps the model identify correct nodes more accurately in graph alignment. These findings confirm that our method can consistently generate high-quality reasoning chains, which in turn lead to significant performance improvements.
Table 5. Quantitative Comparison of Different Reasoning Quality.
Method ExplaGraphs SceneGraphs WebQSP
[…]
$\text{Reasoning}_{\text{Align-GRAG}}$ 0.8992 0.8804 0.5445 0.7626 0.5700
$\text{Reasoning}_{\text{Naive}}$ 0.8592 0.8404 0.5209 0.7444 0.5409
Table 6. The case for summarizing the reasoning chain. The text from SUMMARIZE_PROMPT, Question, Answer, and Graph DataBase are concatenated as input to generate the Response from LLM. We can see that LLM’s summarization accurately identified middle and critical node in graph, ‘J.K. Rowling’, for answer.
SUMMARIZE_PROMPT: You are a helpful assistant responsible for generating a comprehensive summary of the data provided below. Given question and answer, and related graph data base. Please concatenate all of these into a single, comprehensive description. The description should logically connect the question to the answer. Make sure to include information collected from all descriptions.
Question: what is the name of the first harry potter novel?
Answer: harry potter and the philosopher’s stone
Graph DataBase:
@@ 5.10.2. Quantitative comparison
Summarized Reasoning Chain: J.K. Rowling wrote the first Harry Potter novel, “Harry Potter and the Philosopher’s Stone.”
5.11. Case study

### v3

@@ 4.4 Ablation Study
Align-GRAG 0.5445 0.7626 0.5700
w/o Graph Alignment 0.5423 0.7578 0.5673
w/o Node Alignment 0.5344 0.7371 0.5339
w/o Both 0.5348 0.7328 0.5216
Random Alignment 0.4617 0.6861 0.4865
Table 3: Ablation study on different alignment strategy.
@@
(b) Align with Rationale
@@
Figure 3: Graph-level Alignment Analysis: The cosine similarity score between graph embeddings and language embeddings (aligned using the aligner module vs. the unaligned setting).
4.5 Evaluation of Graph-level Alignment
In this section, we evaluate whether the aligner can effectively bridge the representation gap? Specifically, we measure cosine similarity between graph embeddings (unaligned vs. aligned via our contrastive loss) and language embeddings of the query, rationale, and textualized graph on the test set; Figure 3 shows a clear upward trend in similarity as alignment proceeds, with aligned embeddings consistently surpassing unaligned ones after sufficient training, indicating that the aligner effectively narrows the graph–language representation gap. Meanwhile, as discussed in Section C.2, the benefit saturates beyond a certain point, since overly strong alignment can distort original graph information and hurt accuracy.
Figure 4: Quantitative Comparison of Different anchor and rationale Quality.
4.6 Quality Analysis
We investigate whether the quality of extracted anchors and rationale directly affects QA performance. We compare Align-GRAG with No-evidence Extraction, where the LLM generates anchors and a rationale chain using only the question (without accessing the retrieved subgraph evidence). As shown in Figure 4, Align-GRAG consistently performs better across datasets and metrics. This gap indicates that evidence-grounded extraction produces more relevant and faithful supervision signals for both node- and graph-level alignment. In contrast, when extraction is detached from graph evidence, the generated chains are more prone to plausible-but-ungrounded reasoning, which weakens pruning and representation alignment, ultimately harming QA.
@@ 4.6 Quality Analysis
No Pruning
Non-Retriever 100626.28 (387.02%) OOM -
BM25 2569.57 (99.72%) 17:24 min 0.4394
Contriever 2593.28 (100.01%) 17:37 min 0.4429
G-Retriever 2576.66 (100.00%) 17:28 min 0.4502
Node Pruning with Different Degree
@@
Table 4: Evaluation of Efficiency with and without node pruning for inference time. $\uparrow$ indicates results better than G-Retriever, while $\downarrow$ indicates worse results. (xx%) represents the percentage of tokens relative to G-Retriever. OOM means out of memory.
4.7 Evaluating the Efficiency of Node Pruning
@@
In this section, we evaluate the efficiency improvements brought by node pruning in Aligner module. In an inference-only setting on WebQSP, we compare retrieval variants (BM25, Contriever, G-Retriever, and w/ Aligner with different seed-node choices). To fully utilize resources and memory, we adjusted the batch size to be as large as possible under different average token conditions. For comparison, the BM25 and Contriever methods retrieve triples and ensure that the token consumption is roughly similar to that of G-Retriever. Table 4 shows that with a moderate number of seed nodes, our pruning preserves core evidence while removing redundant context—achieves performance comparable to G-Retriever while only utilizing 27.11% tokens, whereas too few seeds can over-prune and hurt performance, and more seeds retain more useful knowledge to outperform G-Retriever in longer-context regimes, indicating that selective pruning improves both efficiency and effectiveness. Finally, under similar token budgets, BM25 and Contriever lag behind, suggesting that graph-aware retrieval better exploits relational connectivity.
5 Related Work
Retrieval-Augmented Generation (RAG)
combines information retrieval to address issues like hallucination and outdated information Gao et al. (2023); Zhou et al. (2025); Xu et al. (2024c). By integrating pre-retrieval (e.g., query rewriting), retrieval, post-retrieval (e.g., reranking), and generation Fan et al. (2024); Fang et al. (2025), RAG Sui et al. (2025); Zamani and Bendersky (2024); Yang et al. (2024); Sudhi et al. (2024); Dong et al. (2024) improves the relevance of retrieved documents and enhances response reliability.
Large Language Models on Graph.
Graphs model relational data in various domains, and GNNs Shi et al. (2020); Veličković et al. (2018); Kipf and Welling (2016) are effective for encoding graph structures. Recent works integrate LLMs with graph learning to enhance graph-based ACL tasks Huang et al. (2024a); Jin et al. (2024); Jiang et al. (2025); Ren et al. (2024). Approaches range from feeding graph tokens into LLMs Tang et al. (2024); Perozzi et al. (2024); Chai et al. (2023) to embedding graph neural layers within transformer architectures Qin et al. (2023); Zhu et al. (2024); Huang et al. (2024b); Wang et al. (2025).
Graph RAG.
Graph RAG incorporates graph-structured knowledge to enhance retrieval and reasoning Baek et al. (2023); Wang et al. (2024); Han et al. (2024); Xu et al. (2024b). It utilizes graph databases Bollacker et al. (2008); Vrandečić and Krötzsch (2014) to retrieve triples or subgraphs Edge et al. (2024); Mavromatis and Karypis (2024); Hu et al. (2025); Wu et al. (2024). Techniques like graph construction from text Guo et al. (2024); Fan et al. (2025) and GNN-based encoding Liu et al. (2025); Jiang et al. (2022); He et al. (2024); Tian et al. (2024) enhance reasoning. Recent approaches leverage LLMs for iterative reasoning over graphs Jiang et al. (2024); LUO et al. ; Sun et al. ; Chen et al. (2024). While Graph RAG emphasizes retrieval, it often lacks robust post-retrieval strategies. Our Align-GRAG introduces a dual alignment mechanism to better link retriever and generator, improving subgraph-query alignment.
6 Conclusion
In this paper, we study GRAG and identify two key obstacles: (i) structure-coupled irrelevant evidence introduced by subgraph expansion and (ii) the structure-reasoning discrepancy between graph encoders and LLM semantics. To address these challenges, we propose Align-GRAG, an anchor-and-rationale guided refinement framework. Experiments on three challenge benchmarks show consistent improvements over 18 strong baselines, demonstrating the effectiveness of the proposed approach. In addition, Align-GRAG exhibits strong generalization across diverse GNN architectures and LLM backbones, indicating robust transferability to different encoder–reasoner pairings. Notably, the aligner’s structure-aware pruning substantially reduces input tokens and inference time while maintaining or even improving accuracy, highlighting Align-GRAG as a practical refinement approach for efficient graph-grounded generation.
Limitations
Despite the promising results demonstrated by Align-GRAG, this work has certain limitations. Due to resource constraints, we were unable to conduct experiments on larger LLMs, leaving the effectiveness of the proposed alignment approach on more powerful models uncertain. Additionally, since our method requires the generation and utilization of graph embeddings, it cannot be directly implemented on closed-source models, which restricts access to internal embedding representations. These limitations highlight potential areas for future exploration, such as validating the scalability of the approach with state-of-the-art LLMs and developing techniques to adapt Align-GRAG for closed-source environments.
Ethical considerations
Our method reduces hallucination by grounding generation in retrieved graph evidence, but it can still produce misleading answers if retrieval returns incomplete/noisy subgraphs or if extracted anchors/rationales are weakly grounded. Since the pipeline relies on LLMs for extraction and assessment, model biases and errors may propagate to both supervision and evaluation; therefore, outputs should be verified before use in high-stakes settings. What is more, we only use LLMs to polish the writing in this work.
Appendix A Graph Retrieval
Existing RAG methodologies are mainly designed for plain text documents or triplets, where information retrieval occurs independently of the graph structure Fan et al. (2024); Zhu et al. (2023); Pan et al. (2024). In the retrieval stage, we first utilize an encoder-only language model (e.g., SentenceBERT Reimers and Gurevych (2020)) to encode textual information, including the query $t_{q}$, text of each node $t_{n}$ and edge $t_{e}$ in the graph, respectively:
$\boldsymbol{q}$ $=\text{SBERT}(t_{q})\in\mathbb{R}^{d},$ (11)
$\boldsymbol{n}$ $=\text{SBERT}(t_{n})\in\mathbb{R}^{d},$ (12)
$\boldsymbol{e}$ $=\text{SBERT}(t_{e})\in\mathbb{R}^{d}$ (13)
$\mathcal{V}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{n}\in\mathcal{V}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{n})\big),$ (14)
$\mathcal{E}_{k}$ $=\operatorname*{arg\,topk}_{\begin{subarray}{c}\boldsymbol{e}\in\mathcal{E}\end{subarray}}\big(\text{sim}(\boldsymbol{q},\boldsymbol{e})\big)$ (15)
This forms top-$k$ entities and relation sets, denoted as $\mathcal{G}_{\text{retriever}}$. Inspired by He et al. (2024), we further leverage the Prize-Collecting Steiner Tree algorithm (PCST) Bienstock et al. (1993) to maintain a controlled graph size.
[…]
$\mathcal{G}_{\text{retriever}}=\underset{\begin{subarray}{c}S\subseteq\mathcal{G},\\ S\text{ is connected}\end{subarray}}{\arg\max}$ $\Bigg(\sum_{n\in\mathcal{V}_{S}}\text{prize}(n)$ (17)
$+\sum_{e\in\mathcal{E}_{S}}\text{prize}(e)$ $-\text{cost}(S)\Bigg).$
where $\mathcal{V}_{S}$ and $\mathcal{E}_{S}$ are the sets of nodes and edges in the subgraph $S$, respectively. The cost of constructing the subgraph is defined as $\text{cost}(S)=|\mathcal{E}|\cdot C_{e}$, where $|\mathcal{E}|$ is the number of edges, and $C_{e}$ is a predefined per-edge cost that serves as a regularization parameter to control the subgraph’s size. In this way, we can obtain a preliminary retrieved subgraph.
Appendix B Experimental Details
Dataset #Training #Validation #Test Task Type
ExplaGraphs 1,659 553 554 Commonsense reasoning
SceneGraphs 59,978 19,997 20,025 Visual scene understanding
WebQSP 2,826 245 1,628 Knowledge-based QA
Table 5: Statistics of GraphQA Benchmark He et al. (2024).
B.1 Datasets and Metrics.
Following G-Retriever He et al. (2024), we conducted experiments on the GraphQA benchmark He et al. (2024), which includes ExplaGraphs (commonsense reasoning), SceneGraphs (scene graph understanding), and WebQSP (knowledge graph reasoning). The dataset statistics are summarized in Table 5, with a data split of train:validation:test = 6:2:2.
B.2 Baselines.
B.2.1 Inference-only.
  • Zero-CoT Wei et al. (2022): enhances zero-shot reasoning by appending the phrase “Let’s think step by step.”, which encourages the model to explicitly generate intermediate reasoning steps.
  • CoT-BAG Wang et al. (2024): adds “Let’s construct a graph with the nodes and edges first.” after providing the textual graph description,
  • KAPING Baek et al. (2023): augments LLMs with knowledge graph facts at the prompt level, enabling more accurate QA without requiring model fine-tuning retrieving relevant graph triples.
  • ToG Sun et al. : integrates LLMs with KGs to support deep and reliable reasoning through interactive exploration of entities and relations.
B.2.2 Raw Fine-tuning.
In this setting, following He et al. (2024), we fine-tune the LLM via parameter-efficient methods, without applying advanced reranking techniques. Two widely used approaches are included:
  • Prompt Tuning Lester et al. (2021): learns a set of continuous prompt embeddings that guide the frozen LLM during inference, enabling lightweight adaptation to specific tasks.
  • LoRA Hu et al. (2021): introduces low-rank adaptation layers into transformer weights, achieving efficient fine-tuning with significantly reduced computational and memory costs.
B.2.3 Reranker-based.
  • GTE (General Textual Embedding) Li et al. (2023): developed by Alibaba DAMO Academy, with two variants (gte-base, 109M parameters; gte-large, 335M parameters), trained on large-scale relevance pairs to improve retrieval accuracy,
  • BGE (BAAI General Embedding) Xiao et al. (2024): cross-encoder models with two variants (bge-reranker-base, 278M parameters; bge-reranker-large, 560M parameters), optimized for retrieval-augmented generation, offering high accuracy at the cost of efficiency,
  • G-RAG and G-RAG-RL Dong et al. (2024): reranker models that leverage GNNs to incorporate document connections and semantic cues from abstract meaning representation graphs for context-aware ranking.
B.2.4 GNN-based.
  • GraphToken Perozzi et al. (2024): encodes graph structures as explicit prompts for LLMs, improving performance on graph reasoning tasks,
  • GNP Tian et al. (2024): a plug-and-play framework that applies a GNN encoder with cross-modality pooling to enrich LLMs with knowledge graph information,
  • G-Retriever He et al. (2024): introduces the GraphQA benchmark and employs retrieval-augmented generation with soft prompting, improving graph-based question answering, comprehension, and reducing hallucinations,
  • GRAG Hu et al. (2025): proposes a divide-and-conquer strategy for efficient textual subgraph retrieval and combines text and graph views within LLMs for graph context-aware generation.
B.3 Implementation Details.
In this section, we further present the Implementation Details. To implement baselines, for GNP Tian et al. (2024), we implemented the Graph Neural Prompting module within our framework, including components such as the GNN encoder, cross-modality pooling, and domain projector. For G-RAG and G-RAG-RL Dong et al. (2024), we adopted their ranking approach, combining cross-entropy loss with pairwise ranking loss. However, since treating documents as nodes was infeasible in our case, we instead treated entities as nodes and employed the same GNN encoder as our method. We implemented GTE Li et al. (2023) and BGE-reranker Xiao et al. (2024) using their official open-source models. Nodes and triples are ranked by query similarity and, as in our method, fed into the LLM for generation. For the $\text{GRAG}_{PCST}$ model Hu et al. (2025), we reproduced its experiments. However, to ensure a fair evaluation, we standardized the retrieval process by using the PCST method for retrieval. This allowed us to directly compare it with their graph encoder approach. Table 6 summarizes the hyperparameter settings for all datasets used in our experiments.
Hyperparameter Value
Number of Seed Nodes 25
Align Degree (Step) 60
Top-$K$ Retrieval 10
Batch Size 8
Max Epochs 20
Learning Rate $1\times 10^{-5}$
GNN Layers 4
GNN Hidden Dim 1024
LoRA Rank 8
LoRA Alpha 16
LoRA Target Modules q_proj, v_proj
Table 6: Hyperparameter settings used in our experiments.
Appendix C Impact of Hyperparameters
@@
C.1 Impact of Seed Nodes
@@
Figure 5: Hyperparameters Analysis of the Number of seed nodes.
C.2 Impact of Align Degree
This section examines how the Align Degree (number of training steps for Aligner module) influences model performance. As shown in Figure 6, we evaluate the Hit@1, F1 and Accuracy metrics on the WebQSP dataset. From the experimental curve of Align Degree, we observe that the Hit@1, F1 and Accuracy metrics peak at around 60 epochs before declining. This indicates that, as training progresses, bridging the representation gap between the graph and language helps the LLM better understand graph data. However, excessive training may lead to overfitting, which disrupts graph information and ultimately causes a drop in performance. This suggests that there is an optimal point in training where the graph alignment is most effective, and training beyond this point can be detrimental.
Figure 6: Hyperparameters Analysis of Align degree.
C.3 Joint Impact of Seed Nodes and Align Degree
To further investigate the interaction between the number of seed nodes ($n_{\text{seed}}$) and the Align Degree (aligner training steps), we report a two-dimensional performance landscape in Figure 7. Each heatmap cell corresponds to one hyperparameter pair $(n_{\text{seed}},\text{Align Degree})$, evaluated on WebQSP using Accuracy, Hit@1, and F1. Overall, the three heatmaps exhibit a highly consistent trend: performance is maximized in a moderate regime, centered around $n_{\text{seed}}\!\approx\!25$ and Align Degree $\approx\!60$. When $n_{\text{seed}}$ is small (e.g., 5), all metrics remain low across nearly all Align Degrees, indicating that insufficient retrieved evidence cannot be compensated by longer aligner training. In contrast, when $n_{\text{seed}}$ becomes large (e.g., 30), the performance degrades, and the degradation is more pronounced under larger Align Degrees, suggesting that excessive retrieved noise may be amplified by over-training the aligner and can eventually harm the graph-language alignment quality. Similarly, for a fixed $n_{\text{seed}}$ in the mid-to-high range, increasing Align Degree improves results initially but starts to decline after the peak region, matching the overfitting pattern observed in the 1D analysis. These results confirm that the best performance requires balancing evidence coverage and noise control, where moderate $n_{\text{seed}}$ ensures adequate reasoning support while the aligner is trained sufficiently (but not excessively) to bridge the representation gap.
Figure 7: Hyperparameter interaction heatmaps on WebQSP.
C.4 Impact of Top K retrieval
@@
Figure 8: Effect of Top K retrieval.
Case Study of Extracted Rationale & Anchors Extraction_Prompt
Question
: what is the name of the first harry potter novel?
Answer: harry potter and the philosopher’s stone  Graph DataBase:
@@
  Extracted Rationale: The question asks for the name of the first harry potter novel, and the graph includes a specific book node labeled harry potter and the philosopher’s stone. The graph’s provided summary states that j. k. rowling wrote the first Harry Potter novel, explicitly naming it as harry potter and the philosopher’s stone. Therefore, using the graph evidence, the first Harry Potter novel is harry potter and the philosopher’s stone, matching the given answer. Extracted Anchors: entity: harry potter and the philosopher’s stone entity: j. k. rowling relation: book.author.works_written
Figure 9: The case for extracting the anchors and rationales. The extracted rationale chain identifies the critical intermediate node j. k. rowling and links it to the answer harry potter and the philosopher’s stone.
Appendix D Case study